A newspaper article reported that people spend a mean of 7.5 hours per day watching TV, with a standard deviation of 1.8 hours. A psychologist would like to conduct interviews with the of the 10 % population who spend the most time watching TV. She assumes that the daily time people spend watching TV is normally distributed. At least how many hours of daily TV watching are necessary for a person to be eligible for the interview? Carry your intermediate computations to at least four decimal places. Round your answer to one decimal place.

I seriously am stuck, anything would help. Thank you
Answer #1

Given that,

mean = = 7.5

standard deviation = = 1.8

Using standard normal table,

P(Z > z) = 10%

= 1 - P(Z < z) = 0.10  

= P(Z < z ) = 1 - 0.10

= P(Z < z ) = 0.90

= P(Z <1.282) = 0.90  

z = 1.282

Using z-score formula  

x = z +

x = 1.282 *1.8+7.5

x = 9.8